# Self-generated electrokinetic flows from active-charged boundary patterns

### v1

@@ Acknowledgements.
The set-up of this numerical scheme entails an axisymmetric cylindrical geometry where a capillary of radius $R=0.1$ $\mu$m and length $L=18$ $\mu$m connects two identical reservoirs as shown in Fig. 5, similar to configurations employed in previous works. The boundary conditions at the capillary surface are the same as those employed in the main text with Eqs. (9) and (10). At the far edges of the capillary-reservoir system of size $W=20$ $\mu$m by $H=1$ $\mu$m, we fix the concentration to the bulk value $c_{\pm}=c_{\infty}=C_{\infty}/2$ and set electric potential to zero $\phi=0$. The regions connecting the reservoirs to the capillary are set to satisfy the condition of zero normal electric field $\nabla\phi\cdot\hat{n}=0$ and also the zero normal flux $\mathbf{J}_{\pm}\cdot\hat{n}=0$. Finally, we impose no-slip boundary conditions for the velocity on all the walls. In COMSOL we conduct a stationary (time-independent) study of this set-up by implementing an interface coupling between the Electrostatics (es) and the Transport of Diluted Species (tds) interfaces provided by the Chemical Reaction Engineering module, in conjunction with the Creeping Flow (spf) interface provided by the Computational Fluid Dynamics module [57]. This scheme which amounts to computing the numerical solutions of the coupled system Eqs. (4)-(8) with the described boundary conditions is executed for a given set of input parameters.
@@
Numerically convergent solutions require appropriate sizes of mesh grid elements in accordance to the system length scales over which our solutions are expected to vary. These scales rely here on parameters such as the Debye length and the flux-charge wavelengths. For instance, larger variations of the electric potential typically occur within the Debye layer near the surface rather than in the bulk far away from it. Thus, near the charged capillary walls a finer mesh is needed with sizes smaller than $\lambda=0.1R$ here. In this respect, we sequentially adjust our mesh sizes inside our capillary domain with a finer mesh of $2\times 10^{-4}R$ near the wall and sparser mesh of $2\times 10^{-2}R$ away from it, and then to larger sizes of up to $0.1H$ in the reservoirs (see Fig. 5). As a result, with these settings COMSOL provides convergent solutions of the electric potential, ion concentrations and fluid velocity for the set-up and parameters employed here.

### v2

@@ Acknowledgements.
The set-up of this numerical scheme entails an axisymmetric cylindrical geometry where a capillary of radius $R$ and length $L$ connects two identical reservoirs as shown in Fig. 7, similar to configurations employed in previous works. The boundary conditions at the capillary surface are the same as those employed in the main text with Eqs. (9) and (10). The regions connecting the reservoirs to the capillary are set to satisfy the condition of zero normal electric field $\nabla\phi\cdot\hat{n}=0$ and also the zero normal flux $\mathbf{J}_{\pm}\cdot\hat{n}=0$. Finally, we impose no-slip and no-penetration boundary conditions for the velocity on all the bounding walls. Finally, at the far edges of the capillary-reservoir system of size $W$ by $H$, we fix the concentration to the bulk value $c_{\pm}=C_{\infty}/2$ and set electric potential to zero $\phi=0$ with the fluid velocity to zero $\mathbf{v}=\mathbf{0}$. In COMSOL we conduct a stationary (time-independent) study of this set-up by implementing an interface coupling between the Electrostatics (es) and the Transport of Diluted Species (tds) interfaces provided by the Chemical Reaction Engineering module, in conjunction with the Creeping Flow (spf) interface provided by the Computational Fluid Dynamics module [66]. This scheme which amounts to computing the numerical solutions of the coupled system Eqs. (4)-(8) with the described boundary conditions is executed for a given set of input parameters.
@@
The system sizes are set according to the capillary radius and wavelength, for instance in Fig. 3 we use $L=180R$, $W=200R$ and $H=10R$. While for the patch model in Fig. 5 we use $L=10$ $\mu$m, $W=30$ $\mu$m and $H=10$ $\mu$m. Numerically convergent solutions require appropriate sizes of mesh grid elements in accordance to the system length scales over which our solutions are expected to vary. These scales rely here on parameters such as the Debye length and the flux-charge wavelengths. For instance, larger variations of the electric potential typically occur within the Debye layer near the surface rather than in the bulk far away from it. Thus, near the charged capillary walls a finer mesh is needed with sizes smaller than $\lambda=0.1R$ here. In this respect, we sequentially adjust our mesh sizes inside our capillary domain with a finer mesh of $2\times 10^{-4}R$ near the wall and sparser mesh of $2\times 10^{-2}R$ away from it, and then to larger sizes of up to $0.1H$ in the reservoirs (see Fig. 7). As a result, with these settings COMSOL provides convergent solutions of the electric potential, ion concentrations and fluid velocity for the set-up and parameters employed here.